OpenStudy (anonymous):

Find the area of a triangle with a length of 12 inches and a with of 10 inches.

OpenStudy (diyadiya):

Area of a triangle or Rectangle? \[Area~of~a~\triangle = \frac{1}{2} \times base \times height\]\[Area~of~a~Rectangle = length \times width \]

OpenStudy (anonymous):

Does it say width or height?

OpenStudy (anonymous):

@Diyadiya Sorry! I meant a rectangle. @monokerous *width ._.

OpenStudy (anonymous):

ok. rectangle is length*width.

OpenStudy (anonymous):

So what would the radius be? @monokerous

OpenStudy (anonymous):

rectangles have no radiuses.

OpenStudy (anonymous):

i'm assuming you meant area? 12*10=?

OpenStudy (anonymous):

I'm all confused,sorry. This is what it says "Find the area of a triangle with a length of 12 inches & a width of 10 inches" I was looking at the wrong question,haha.

OpenStudy (anonymous):

12*10=? Solve the equation. That will be your answer.